Lisa is conducting an expeeriment of tossing two ordinary dice and adding the two numbes. Determine the probability of her obtaining:
a) A sum of 5
b) A sum greater than or equal to 8
c) A sum less than 6

Answer by stanbon(75887)

Lisa is conducting an expeeriment of tossing two ordinary dice and adding the two numbes. Determine the probability of her obtaining:
Procedure: You should draw the sample space as a 6x6 square
with the rows and columns headed 1,2,3,....6 and the square
inside the sum of the row and column numbers.
--------------
a) A sum of 5
count the ways you can get at a 5: (1,4), (2,3), (3,2), (4,1)
P(sum of 5) = 4/36 = 1/9
----------------
b) A sum greater than or equal to 8
count and divide by 36
------------
c) A sum less than 6
count and divide by 36
====================
Cheers,
Stan H.